The difference between two numbers is 27. 4 times the larger number is 6 times the smaller number. Find the two numbers.
Answer by london maths tutor(243)

Let's the large number be L and small number be s
L - s = 27
4L = 6s
s = 4L/6
L - (4L/6) = 27
6L - 4L = 27 * 6
2L = 162
L = 81
81 - s = 27
s = 81 - 27
s = 54
The two numbers are 81 and 54